Status Report

ISS On-orbit Status 23 Jan 2003

By SpaceRef Editor
January 23, 2003
Filed under , ,
ISS On-orbit Status 23 Jan 2003

All ISS systems continue to function nominally, except as noted previously
or below.

The crew awoke to congratulations on yesterday’s PAO/educational event
with NASA GRC (Glenn Research Center) and was informed that the students
really enjoyed the crew’s answers and the video, especially when props
were used.

Today’s Robotics ops were successfully completed. Starting at about
3:15am, CDR Ken Bowersox and FE-2/SO Don Pettit maneuvered the Canadarm2/SSRMS
(space station remote manipulator system) into the proper position to
view the S1 truss radiator clearance which had caused some concern during
the checkout of its TRRJ (thermal radiator rotary joint). [Last night
before sleep time, the crew connected the Lab RWS DCP (robotics workstation
display and control panel) bypass power cable, and MCC-H had the MSS
(mobile service system) powered up this morning by 2:00am. Robotics activities
began with the usual hand controller calibration and setup for the first
maneuver, working "in the blind" with the DOUG (dynamic operations
ubiquitous graphics). They then "walked" the robotarm from
its the current base, the Lab PDGF (power and data grapple fixture),
to MBS (mobile base system) PDGF2 (new base), by using MBS PDGF1 as a "way
station" and the TE (tip elbow) camera to check on clearance to
the S-band antenna. After subsequent release of MBS PDGF1, a quick grapple/release
of MBS PDGF3 as an SSRMS digitals accuracy characterization exercise
was also part of the move to S1 clearance viewing position.]

FE-1 Nikolai Budarin meanwhile performed the second experiment run of
the Russian Plasma Crystal-3 (PK-3) payload, setting up the hardware
and video recording. The turbopump was to be deactivated again in the
evening. [During the day, Budarin monitored the experiment, which today
searched for mixture modes without voids in the center of the plasma
(charged fine particles in the vacuum chamber) at various discharge power
and various pressures, and studies residual charge of small particles
after discharge is removed (relaxation) at various pressures, wave dispersion
in mixtures, and solitary wave dispersion in mixtures at reduced pressure.
The first part was then repeated.]

Budarin also conducted his regular daily inspection of the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment. He
later copied its photo/data files to a floppy disk for transfer to the
Russian Laptop 3 and subsequent downlink via Regul-Packet.

Bowersox prepared the data from yesterday’s FOOT foot/ground reaction
forces during space flight) session and transferred the file for subsequent
downlink to MCC-H, which he finished by about 12:55pm.

Budarin and Pettit completed another session with the Russian MedOps
cardio experiment MO-1 (study of the bioelectric activity of the heart
at rest), with Don assisting Nikolai as CMO (crew medical officer), and
tagging up with a ground specialist at TsUP (MCC-M).

Budarin also continued his periodic maintenance work on the RS (Russian
segment) air circulation system, today working in the "Pirs" DC-1
docking module to change out its two dust filters (PF1-2) and cleaning
the mesh screens of the ventilator fans (V1-2).

Later, Nikolai removed three devices from the BD database monoblock
(TA985M) which he had replaced in the BITS2-12 on-board telemetry measurement
system with a spare unit. [The three modules (TA750, TA837B, TA746) are
to remain on orbit as spares, whereas the TA985M is to be returned to
the ground.]

Pettit performed the regular daily routine task of IMS inventory delta
file preparation for automated downlink, while Budarin completed the
daily routine maintenance/inspection of the SOSh life support system
(incl. ASU toilet subsystem).

Don Pettit also collected the data of the FMK (formaldehyde monitoring
kit) monitors deployed yesterday in the Lab and Service Module (SM).

On MCC-M go-ahead, the ISS cabin atmosphere was pressurized with fresh
oxygen (O2) from the Progress 9P O2 tank.

As a task list item, Sox and Don at an earlier time filled out their
weekly FFQ (food frequency questionnaire), a special software log on
the MEC (medical equipment computer) to track nutritional intake of the
two crewmembers.

Via live TV downlink from the SM, the crew addressed participants of
the 31st Russian National Space Competition in Korolev near Moscow during
today’s closing ceremony. The two-day finale of the student competition
was dedicated to the 40th anniversary of the first group space flight
(Vostok-5/Vostok-6). [The competition has become a tradition, having
been held annually since 1971 to "encourage children’s creative
abilities, to support talented youth, and help educate Russia’s intellectual
elite". Sponsors are the Soyuz Russian National Youth Aerospace
Association, the Ministry of Education of the Russian Federation, Rosaviakosmos,
the Moscow City Government, the Moscow Region Government Administration,
the Khrunichev Space Center, and RSC Energia in Korolev. The finale was
attended by over 250 participants from 42 Russian cities, (including
Baikonur), as well as from Belarus and Ukraine. The jury consisted of
cosmonauts, scientists, rocket-space technology engineers, and representatives
from the major aerospace university departments in Moscow.]

At 9:38am, Science Officer Dr. Don Pettit participated in an amateur
(ham) radio chat with Grade 3-5 students at Eugene Field School in Park
Ridge, Illinois. [More than 600 students attend this school, which counts
as its most famous alumni U.S. Senator Hilary Clinton and actor Harrison

POC (Payload Operations Center, Huntsville) is still trying to determine
why Express Rack 2 (ER2) is not communicating with its laptop. The troubleshooting
activity has today been added to the crew’s "job jar" task

When Pettit inspected the coils of the InSPACE (investigating the structure
of paramagnetic aggregates from colloidal emulsions) experiment to determine
their viability in preparation for Progress 10P arrival and MSG (microgravity
science glovebox) activation, he found that one of the capillary tubes
appears to have been damaged and the contents leaked out. [This is no
impact to InSPACE science objectives as there are two sets of each of
the three coils onboard and the spare coil is intact. Pettit found that
the remaining coil assemblies showed some settling of the fluid and that
by shaking the coils he was able to evenly disperse the media.]

On 1/17, after multiple attempts, Pettit was unable to successfully
complete the TOCA (total organic carbon analyzer) water sampler reagent
mixing. Analysis now shows that TOCA can probably not be repaired on-orbit.
[TOCA is a water quality monitoring instrument for determining concentrations
of total organic carbon, total inorganic carbon, total carbon, pH value
and conductivity in water. The failed TOCA has been manifested for return
on ULF1. Real-time monitoring of potable water is no longer possible
for the remainder of Increment 6, but post-flight sampling with the TOCA
is still planned.]

Yesterday’s water venting from the Lab condensate tank and two CWCs
(contingency water containers) required a total propellant expenditure
of about 20 kg for attitude hold in -YVV/Barbecue (inverted with zenith
toward Earth), well within the prop limit of 25 kg set beforehand by

Today’s targets for the CEO (crew earth observations program), limited
because daylight passes are currently mainly over the oceanic hemisphere,
were Cape Town, South Africa (nadir pass; ESC [electronic still camera]),
Lake Eyre, Australia (this non-outlet basin in east-central Australia
[large white-floored lake bed]) responds to various climatic controls.
A major control is the El Nino periodicity [presently underway], when
the surrounding basin generally receives less rainfall), and Buenos
Aires, Argentina (nadir and a touch right; ESC).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:25pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber
is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control
subsystem) is off. MCA (major constituents analyzer) is operating.
BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2
in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off.

SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C)
— 25.9; ppO2 (mmHg) — 160.9; ppCO2 (mmHg) — 2.3 (suspect).
SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C)
— 19.8.
FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 23.7.
Node: Pressure (mmHg) — 740.07; temperature (deg C) — 24.1 (shell);
ppO2 (mmHg) — 162.4; ppCO2 (mmHg) — 4.4.
U.S. Lab: Pressure (mmHg) — 741.99; temperature (deg C) — 22.7; ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 742.09, temperature (deg
C) — 22.0; shell heater temp (deg C) — 24.3, ppO2 (mmHg) — n/a; ppCO2
(mmHg) — n/a.
PMA-1: Shell heater temp (deg C) — 24.3
PMA-2: Shell heater temp (deg C) — 11.5
(n/a = data not available)

Propulsion System (PS): Total propellant load available (SM + FGB +
Progress) — 3669 kg (8089 lb) [as of 1/23/03].

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B both in Autotrack (sun-following) mode.
SM batteries (as of 3am): Battery #7 is off line (failed); battery #4
is off line (ROM/capacity restoration mode); all other batteries (6)
are in "Partial Charge" mode.
FGB batteries (3am): Battery #3 is offline (failed); battery #5 is in "Cycle
mode"; all other batteries (4) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-1 is On (primary), EXT-2 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane
1 is down (as of 11/14).
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line (CMG-1 failed).
State vector — U.S. SIGI-1
Attitude — U.S. SIGI-1
Angular rates — U.S. RGA-1

Flight Attitude:
XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw:
-179.5 deg, pitch: -8.8 deg., roll: 0 deg]), CMG/Thruster Assist Momentum
Solar Beta Angle: 31.6 deg (magnitude decreasing).

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operating.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both
MBS: KA power on both strings.
MT: latched at WS4, with KA power.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning,7:23am EST [= epoch]):
Mean altitude — 388.7 km
Apogee — 392.7 km
Perigee — 384.8 km
Period — 92.33 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005844
Orbits per 24-hr. day — 15.60
Altitude loss — 150 m (mean) in last 24 hours
Solar Beta Angle — 31.6 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 23843

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.